# Disclosure of AI-Generated News Increases Engagement but Does Not Reduce Aversion, Despite Positive Quality Ratings

### v1

@@ Abstract
The advancement of artificial intelligence (AI) has led to its application in many areas, including journalism. One key issue is the public’s perception of AI-generated content. This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future. We employed a between-subjects survey experiment with 599 participants from the German-speaking part of Switzerland, who evaluated the credibility, readability, and expertise of news articles. These articles were either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely generated by AI (AI-generated group). Our results indicate that all news articles, regardless of whether they were written by journalists or AI, were perceived to be of equal quality. When participants in the treatment groups were subsequently made aware of AI’s involvement in generating the articles, they expressed a higher willingness to engage with (i.e., continue reading) the articles than participants in the control group. However, they were not more willing to read AI-generated news in the future. These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could attract more immediate engagement with their content, at least in the short term.
@@ 1 Introduction
The landscape of journalism is evolving significantly with the advancement of artificial intelligence (AI). Traditionally, journalists have used computers primarily for research and data analysis, but this role is rapidly expanding. Modern AI, particularly generative AI, now matches or even surpasses human capabilities in tasks like text composition. This has led to the emergence of “automated” or “robot” journalism (Carlson, 2018; Clerwall, 2014; Graefe et al., 2018), a key aspect of the broader computational journalism trend that underscores the increasing role of computation and data in news production (Anderson, 2013; Cohen, Hamilton and Turner, 2011; Lewis and Westlund, 2015).
AI applications, especially in natural language generation, have become integral in practical journalism. These applications include news writing for media organizations such as the Associated Press and Forbes, summarizing scientific data via platforms like the Open Research Knowledge Graph, and narrative writing using tools like ChatGPT-4. These advancements enable AI to produce content that is virtually indistinguishable from human writing, demonstrating AI’s potential in various writing-intensive domains.
The integration of AI-generated content in journalism presents several implications. On one hand, it offers opportunities for faster, multilingual, and expanded content production with potentially fewer errors, which could enhance news quality and help combat misinformation (Graefe et al., 2018). AI can also handle routine reporting tasks, allowing journalists to focus on investigative stories. On the other hand, however, there are concerns about job losses in the journalism sector and doubts regarding algorithms’ ability to fulfill the “watchdog” role traditionally associated with journalists (Strömbäck, 2005; Latar, 2015).
@@
A crucial issue is the public’s perception of AI-assisted journalism. Understanding attitudes towards AI-generated journalism is essential as it may directly impact the public’s political knowledge. Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry. A recent study revealed that only 29% of Swiss respondents would read fully AI-generated news, 55% would read news if AI assisted in its creation, and 84% would prefer news created without AI involvement (Vogler et al., 2023). This highlights skepticism towards AI-generated news amongst the Swiss population, consistent with surveys from other countries (Fletcher and Nielsen, 2024). Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations. Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
@@
In this preregistered study, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future. Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise. When participants in the treatment groups were subsequently informed about AI’s role in generating the articles, they expressed a higher willingness to read the full news article than participants in the control group, who saw human-written articles. This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
@@
These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is of lower quality. Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement. Future work should delve deeper into understanding the underlying mechanisms driving both results, exploring factors such as the novelty of AI involvement, trust in AI, and the psychological effects of transparency. Additionally, research should investigate long-term impacts on reader behavior and trust to develop strategies for effectively integrating AI in journalism, ensuring that it complements rather than undermines the journalistic process.
@@ 2 Previous Research
Several strands of literature have explored the perception of human versus AI-generated news. The first strand involves research on public opinions about human versus computer-generated news, relying on respondents’ prior experiences or perceptions (e.g., Vogler et al., 2023; Chuan, Tsai and Cho, 2019). A recent survey conducted across six countries (Argentina, Denmark, France, Japan, the UK, and the USA) revealed that public opinion on the use of generative AI in journalism is mixed (Fletcher and Nielsen, 2024). While people expect AI-produced news to be cheaper and more up-to-date, they also anticipate it to be less trustworthy and transparent. Most respondents are wary of AI-generated news, especially on hard topics like politics and international affairs, but show more acceptance for soft news like fashion and sports. A large majority favors disclosure when AI is used, but opinions differ on which specific uses should be labeled. Uncertainty remains, with many respondents unsure or neutral about AI’s role in journalism. A limitation of these kinds of opinion-based studies is that they could reflect respondents’ preconceptions or lack of familiarity with AI-generated texts, which may not align with actual behaviors when confronted with AI news.
The second strand consists of surveys where respondents directly evaluate news stories labeled as either “human-written” or “computer-written.” For instance, Haim and Graefe (2017) found that respondents rated human-written news higher in readability but lower in credibility compared to AI-generated news. However, when participants were only given a single article without information on its origin, the differences in their evaluations were minimal, indicating that the perception of the author (human or AI) plays as significant a role in the assessment as the content itself.
The third strand focuses on experiments where participants are presented with identical articles but with different stated authors, to measure the effect of the source. For example, Van der Kaa and Krahmer (2014) found that while news consumers rated the credibility of both human and computer-generated articles equally, journalists perceived human-labeled articles as significantly more trustworthy. This effect varied with the topic, as sports articles were rated as less trustworthy than financial ones. Jung et al. (2017) used a $2\times 22 × 2$ design to explore perceptions of article quality based on authorship labels, finding that both journalists and the public rated AI-written articles higher when they were correctly labeled as such. Another experiment by Tandoc Jr, Yao and Wu (2020) varied the attributed author of a human-written text, finding no significant difference in perceived credibility across different labels. Additionally, Altay and Gilardi (2023) explored how labeling news headlines as “AI-generated” affects accuracy ratings and sharing intentions, finding that such labels decreased perceived accuracy and sharing intent, even when the text was human-written.
The fourth strand investigates the effect of the message itself by providing texts written by different authors without stating the source, thereby reducing source-based biases. Clerwall (2014) found that computer-written articles were perceived as more credible, whereas human-written articles scored higher in readability. This approach highlights the challenge of distinguishing between human and AI-generated content. Kim et al. (2020) extended this analysis across multiple mediums (text, audio, video), finding significant differences favoring human-generated content only in audio and video formats. In a follow-up study, Kim, Xu and Merrill Jr (2022) discovered that human newscasters were perceived as more credible than AI newscasters.
Some studies merge the third and fourth strands by measuring both source and message effects. For instance, Graefe et al. (2018) examined the effects of declared source, actual article content, and topic, finding that while declared source effects generally favored human-written articles, computer-written articles scored higher in credibility and expertise in fact-based topics. Wölker and Powell (2021) explored European readers’ perceptions of automated journalism, finding that automated and human content were perceived as equally credible.
Finally, a meta-analysis incorporating results from studies across all four strands found no significant difference in credibility between human-written and AI-generated news, with human-written texts only having a slight advantage in perceived quality and a notable lead in readability (Graefe and Bohlken, 2020). However, all studies in this analysis were conducted before 2020, and the rapid advancements in generative AI may have altered these dynamics.
@@
Overall, existing research suggests that while AI-generated news can be perceived as being of equal quality to human-generated news, knowing that an article was created by AI can negatively influence readers’ perceptions. This implies that preconceived notions may lead to distorted preferences for human-generated news. However, previous studies have not separated quality evaluations from the willingness to read AI-generated news, limiting our understanding of public attitudes in this area. Moreover, recent innovations in AI text generation may have made some of the existing findings outdated, underscoring the need for additional research.
@@ 3 Research Questions and Hypotheses
By answering these two research questions, we extend the literature in important ways. First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices. Second, we make sure that respondents’ views on AI are not asked in the abstract, which may reflect inaccurate preconceptions, but are rooted in specific examples we provide. Third, we separate quality evaluations from willingness to read AI-generated news by disclosing the degree of AI involvement after participants evaluated the articles’ quality, which ensures that awareness of AI’s role does not affect quality ratings.
The first set of hypotheses focus on quality assessments of news articles with different degrees of AI involvement. Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality. However, we acknowledge that recent advances in generative AI may challenge these expectations.
@@
Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality.
Graefe et al. (2018, 604) found that “news consumers get more pleasure out of reading human-written as opposed to computer-written content”. Hence, we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
@@ 4 Experimental Design
To explore the public’s perceptions of AI-generated and AI-assisted news articles compared to traditional human-generated articles, we conducted an online between-subjects survey experiment. We recruited 599 participants from the age of 18 from the German-speaking part of Switzerland through the survey company Bilendi, using quotas based on age and gender to ensure balance across these dimensions. Sample size was determined by a power analysis, assuming an effect size of Cohen’s d = 0.3 for the primary outcomes and d = 0.15 for the secondary outcomes.
First, participants completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation. Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
@@ 5 Results
Table 2 provides a detailed overview of the regression results for the second set of outcomes. They reinforce the findings from Figure 2, indicating a higher willingness to continue reading in these groups. For future reading willingness, the coefficients for both the AI-assisted and AI-generated groups are not significant, corroborating the findings from Figure 2 that there is no significant difference in future reading preferences between the different groups. We therefore reject of hypotheses H9 and H10, which predicted higher willingness to read AI-generated news in the future. These results suggest that knowledge of AI involvement and exposure to it do not significantly influence future reading preferences.

### v2

@@ Abstract
The advancement of artificial intelligence (AI) has led to its application in many areas, including news media. The integration of AI in journalism presents both opportunities and risks for democracy, making it crucial to understand public reception of and engagement with AI-generated news, as it may directly influence political knowledge and trust. This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future. We employed a between-subjects survey experiment with 599 participants from the German-speaking part of Switzerland, who evaluated the credibility, readability, and expertise of news articles. These articles were either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely generated by AI (AI-generated group). Our results indicate that all news articles, regardless of whether they were written by journalists or AI, were perceived to be of equal quality. When participants in the treatment groups were subsequently made aware of AI’s involvement in generating the articles, they expressed a higher willingness to engage with (i.e., continue reading) the articles than participants in the control group. However, they were not more willing to read AI-generated news in the future. These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could attract more immediate engagement with their content, at least in the short term.
@@ 1 Introduction
The news media landscape is evolving significantly with the advancement of artificial intelligence (AI). AI applications, especially in natural language generation, have become integral to the production of news (Cools and Diakopoulos, 2024; Simon, 2024b, a). The integration of AI in the news presents both opportunities and risks for democracy (Arguedas and Simon, 2024). On the one hand, AI has the potential to make news more accessible and tailored to diverse audiences, supporting informed public debate. On the other hand, however, challenges arise around transparency, bias, and trust, as AI models often reflect the dominant worldviews of their creators, potentially narrowing the range of perspectives in public discourse.
@@
A specific issue is the public’s perception of AI-produced news. Understanding attitudes towards AI-generated news is essential as it may directly impact the public’s political knowledge and trust. Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry. A recent study revealed that only 29% of Swiss respondents would read fully AI-generated news, 55% would read news if AI assisted in its creation, and 84% would prefer news created without AI involvement (Vogler et al., 2023). This highlights skepticism towards AI-generated news amongst the Swiss population, consistent with surveys from other countries (Fletcher and Nielsen, 2024). Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations. Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
@@
In this preregistered study (see the Appendix for an anonymized version of the pre-registration), conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future. Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise. When participants in the treatment groups were subsequently informed about AI’s role in generating the articles, they expressed a higher willingness to read the full news article than participants in the control group, who saw human-written articles. This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
@@
These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is perceived to be of lower quality. Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement. Future work should delve deeper into understanding the underlying mechanisms driving both results, exploring factors such as the novelty of AI involvement, trust in AI, and the psychological effects of transparency. Additionally, research should investigate long-term impacts on reader behavior and trust to develop strategies for effectively integrating AI in journalism, ensuring that it complements rather than undermines the journalistic process.
@@ 2 Previous Research
AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Diakopoulos, 2024). Structured data-based formats like sports and financial reporting are particularly amenable to AI-generated content due to their formulaic nature. AI can assist with data-heavy tasks, enabling journalists to focus on responsibilities that demand human judgment, creativity, and critical thinking, such as investigative reporting and nuanced writing (Clerwall, 2014; Haim and Graefe, 2017). These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms (Simon, 2024b).
Research on public perceptions of news authored by AI reveals a mixed picture. While earlier studies demonstrated preferences for human-written news (Graefe et al., 2018; Graefe and Bohlken, 2020), more recent work found few notable differences in perceived quality, expertise, credibility, and readability between human-written and AI-generated content (Wölker and Powell, 2021; Kim et al., 2020; Kim, Xu and Merrill Jr, 2022). However, AI-produced content tend to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020). People tend to view AI-produced news as less trustworthy and transparent, and less worth paying for over human-written news (Fletcher and Nielsen, 2024). Most support labeling AI involvement, especially for core tasks like writing, with more acceptance in soft news topics than in hard news. However, public perceptions of AI-generated news often depend on the specific ways AI is implemented and whether or how its use is disclosed (Altay and Gilardi, 2024). AI can be viewed as more objective and less biased than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024). However, perceptions of algorithmic neutrality can be misleading, as AI systems trained on biased data may inadvertently perpetuate or amplify existing biases.
While transparency is generally seen as beneficial for accountability and audience trust (Wang and Ophir, 2024), simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024). People tend to be skeptical of headlines labeled as AI-generated, regardless of whether they are genuinely AI-produced or actually created by humans, partly because they tend to assume an extreme, possibly unrealistic degree of AI involvement (Altay and Gilardi, 2024). Without careful management, transparency can lead to negative biases against AI, reducing credibility, particularly among technology-skeptical audiences (Wang and Huang, 2024; Wang and Ophir, 2024).
AI integration in journalism should be conducted responsibly (Cools and Diakopoulos, 2024). Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations”, and safeguarding journalistic autonomy and core values (Cools and Diakopoulos, 2024). News organizations are encouraged to create structures to experiment with, assess, and refine best practices for incorporating AI tools across all stages of news production, from information gathering to editing and distribution (Araujo et al., 2023; Cools and Diakopoulos, 2024). A collaborative model where AI supports rather than replaces journalists, in conjunction with more effective frameworks for sharing information on AI across every level of news organizations, is generally considered best practice (Dodds et al., 2024).
The integration of AI in journalism presents both opportunities and risks for democracy (Arguedas and Simon, 2024). AI offers the promise of enhancing accessibility and personalization in news delivery, fostering more inclusive and informed discussions. Yet, concerns persist regarding issues like transparency, bias, and trust, as these systems often mirror the perspectives and biases of their developers, potentially limiting diversity in the viewpoints presented to the public.
@@
In sum, existing research suggests that while AI-generated news can be perceived as being of equal quality to human-generated news, knowing that an article was created by AI can negatively influence readers’ perceptions. This implies that preconceived notions may lead to distorted preferences for human-generated news. However, previous studies have not separated quality evaluations from the willingness to read AI-generated news, limiting our understanding of public attitudes in this area. Moreover, recent advancements in AI text generation suggest that some pre-ChatGPT findings may no longer fully apply, emphasizing the importance of continuous, additional research to keep pace with these developments.
@@ 3 Research Questions and Hypotheses
By answering these two research questions, we extend the literature in three important ways. First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices. Second, we make sure that respondents’ views on AI are not asked in the abstract, which may reflect inaccurate preconceptions, but are rooted in specific examples we provide. Third, we separate quality evaluations from willingness to read AI-generated news by disclosing the degree of AI involvement after participants evaluated the articles’ quality, which ensures that awareness of AI’s role does not affect quality ratings.
The first set of hypotheses focus on quality assessments of news articles with different degrees of AI involvement. Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality. However, we acknowledge that more recent studies may challenge these expectations (Wölker and Powell, 2021).
@@
Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality. Given that AI-generated content is often met with distrust, particularly when it involves political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020), we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
@@ 4 Experimental Design
To explore the public’s perceptions of AI-generated and AI-assisted news articles compared to traditional human-generated articles, we conducted an online between-subjects survey experiment (the preregistration containing full study details can be found in the Appendix). We recruited 599 participants from the age of 18 from the German-speaking part of Switzerland through the survey company Bilendi, using quotas based on age and gender to ensure balance across these dimensions. The study was conducted in May 2024. Sample size was determined by a power analysis, assuming an effect size of Cohen’s d = 0.3 for the primary outcomes and d = 0.15 for the secondary outcomes.
After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation. Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
@@ 5 Results
Table 2 provides a detailed overview of the regression results for the second set of outcomes. They reinforce the findings from Figure 2, indicating a higher willingness to continue reading in the AI-assisted (p<0.01) and AI-generated (p<0.01) groups. For future reading willingness, the coefficients for both the AI-assisted and AI-generated groups are not significant, corroborating the findings from Figure 2 that there is no significant difference in future reading preferences between the different groups. We therefore reject of hypotheses H9 and H10, which predicted higher willingness to read AI-generated news in the future. These results suggest that knowledge of AI involvement and exposure to it do not significantly influence future reading preferences.